OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, -1) and (-1, -1)?

OpenStudy (anonymous):

zero since the y coordinates are the same

OpenStudy (karatechopper):

0/4= 0 slope

OpenStudy (anonymous):

o

OpenStudy (anonymous):

slope formula (y2-y1) -------- = m (x2-x1) notice that when they multiply both sides of that by (x2-x1) they get something like this: point slope formula: y-y1 = m(x-x1)

OpenStudy (anonymous):

zero